OpenStudy (anonymous):

Solve the proportion: x/2 = 5/15 x =

OpenStudy (anonymous):

the square root of 2/3

OpenStudy (anonymous):

\[\frac{ x }{ 2 }=\frac{ 5 }{15 }\] Cross multiply: \[15x=10\] Dived both sides byy 15 and you'll get: \[\frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

*Divide *by

OpenStudy (anonymous):

@MommyHatinMath

OpenStudy (anonymous):

how do you get 10

OpenStudy (anonymous):

just kidding 5x2 is 10